I am stuck on this problem, please help!
Complete the ordered pair so that is satisfies the given equation.
a=b^2+4b+5 (-2, ??) (??,2)
Thanks in advance for your help!!

Answer by Earlsdon(6294)

Let's take the first given ordered pair, (-2, ?) which you could write in more familiar terms (x, y) or (-2, y) and rewrite the equation in more familiar notation:
y = x^2 + 4x + 5 Substitute x = -2 and solve for y:
y = (-2)^2 + 4(-2) + 5
y = 4 - 8 + 5
y = 1
The first ordered pair is (-2, 1)
Now the second given ordered pair, (x, 2):
y = x^2 + 4x + 5 Substitute y = 2 and solve for x.
2 = x^2 + 4x + 5 Subtract 2 from both sides.
x^2 + 4x + 3 = 0 Factor.
(x + 1)(x + 3) = 0 Apply the zero products principle.
x + 1 = 0; x = -1 or x + 3 = 0; x = -3
The second ordered pair is: (-1, 2) or (-3, 2)